OpenStudy (anonymous):

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? -x + 2y = -6 3x + y = 8 Answer choices: -2y + 6 2y - 6 2y + 6 -2y - 6

OpenStudy (anonymous):

the 3rd one...

OpenStudy (anonymous):

Can you show me how?.. I tried myself but I struggled with isolating the variable @Orion1213

OpenStudy (anonymous):

\[-x+2y=-6\]we do apply transposition on both sides... by adding x+6 on both sides of the equation... \[x+6-x+2y=-6+x+6\] \[2y+6=x\] \[x=2y+6\]

OpenStudy (anonymous):

... that is to equate x in terms of the remaining terms of variables and constants...

OpenStudy (anonymous):

... got the idea? @Glorenda49

OpenStudy (anonymous):

Yes I understand completely. thank you @Orion1213